# Coding for Gaussian Two-Way Channels: Linear and Learning-Based Approaches

### v1

@@ V Computational Complexity of Linear and RNN-Based Coding
In this section, we analyze the computational complexity of our proposed linear and RNN-based coding schemes. We first look into the case of linear coding. In Algorithm 2, we solve for the encoding schemes, i.e., ${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, and ${\bf F}_{2}$, given the fixed number of channel uses, $N$, and the noise variances, $\sigma_{1}^{2}$ and $\sigma_{2}^{2}$. Subsequently, we obtain the decoding (combining) schemes, ${\bf w}_{1}$ and ${\bf w}_{2}$, by (22) as functions of the encoding schemes. Once the encoding/decoding schemes are determined, they are used for encoding via (10)-(11) and decoding described in Sec. III-B without being recalculated. Specifically, the complexity of encoding at User 1 is $\mathcal{O}(N^{2})$ from (10), where ${\bf F}_{1}{\bf F}_{2}$ can be pre-calculated, while User 2 has the complexity of $\mathcal{O}(N)$ in (11), where the computation of ${\bf F}_{2}{\bf y}_{2}$ has $\mathcal{O}(N)$ complexity since ${\bf F}_{2}$ is in the form of (30). For decoding, User 1 has complexity of $\mathcal{O}(N^{2})$, where the decoding process at User 1 consists of (i) $\tilde{\bf y}_{1}=({\bf I}+{\bf F}_{2}{\bf F}_{1})^{-1}({\bf y}_{1}-{\bf F}_{2}{\bf g}_{1}m_{1})$ and (ii) ${\hat{m}}_{2}={\bf w}_{2}^{\top}\tilde{\bf y}_{1}$, described above (19). Here, $({\bf I}+{\bf F}_{2}{\bf F}_{1})^{-1}$ and ${\bf F}_{2}{\bf g}_{1}$ can be pre-computed. On the other hand, User 2 has $\mathcal{O}(N)$ complexity, where the decoding process at User 2 is composed of (i) $\tilde{\bf y}_{2}={\bf y}_{2}-{\bf F}_{1}{\bf g}_{2}m_{2}$ and (ii) ${\hat{m}}_{1}={\bf w}_{1}^{\top}\tilde{\bf y}_{2}$, described above (21). Here, ${\bf F}_{1}{\bf g}_{2}$ can be pre-computed.
@@
The encoding and decoding complexity at the users for linear coding and RNN-based coding are summarized in Tables II and II. Linear coding has a low coding complexity due to the simplified system model adopting linear operation for encoding and decoding in (10)-(11) and (22). While the RNN-based coding causes higher encoding/decoding complexity, it benefits from higher degrees of flexibility thanks to the non-linearity provided by deep neural networks, leading to better error performance under many practical noise scenarios (shown in Sec. VI).
@@ VI Experimental Results
In this section, we investigate both our linear and learning-based coding methodologies through numerical experiments. We first describe the simulation setup (Sec. VI-A) and validate our coding methodologies in a short block-length regime (Sec. VI-B). In addition, we discuss the enhancement of our linear coding scheme by thoroughly investigating its solution behavior (Sec. VI-C). We also show the effectiveness of our coding methodologies in a medium/long block-length regime (Sec. VI-D). Then, we analyze the benefits of two-way coding (Sec. VI-E), and show that the best coding strategy is different in SNR regimes (Sec. VI-F). Moreover, we investigate the error performance by varying the coding rates (Sec. VI-G) and provide information theoretic insights on power distribution of both linear and RNN-based coding (Sec. VI-H). Finally, we analyze the block-length gain of our RNN-based coding (Sec. VI-I).
@@ VI-A Simulation Parameters
For RNN-based coding, the number of training data samples that we used is $J=10^{7}$, the batch size is $N_{\text{batch}}=2.5\times 10^{4}$, and the number of epochs is $N_{\text{epoch}}=100$. We use Adam optimizer [26] and a decaying learning rate, where the initial rate is 0.01 and the decaying ratio is $\gamma=0.95$ applied for every epoch. We also use gradient clipping for training, where the gradients are clipped when the norm of gradients is larger than 1. We adopt two layers of uni-directional GRUs at the encoder and two layers of bi-directional GRUs at the decoder for each user, with $N_{\text{neurons}}=N_{i}^{(\text{enc})}=N_{i}^{(\text{dec})}=50$ neurons at each GRU. We initialize each neuron in a GRU with the uniform distribution $U(-\zeta,\zeta)$, where $\zeta=1/\sqrt{N_{\text{neurons}}}$, and all the power weights and attention weights to 1. We train our neural network model under particular noise scenarios, i.e., $\text{SNR}_{1}^{\text{ch}}$ and $\text{SNR}_{2}^{\text{ch}}$, and conduct inference under the same noise scenario. For inference, we consider $10^{10}$ data tuples to calculate BER and BLER.
@@ VI-B Short Block-Length
We first investigate the performance of our coding strategies in a short block-length regime. We consider the same block-length of $L_{1}=L_{2}=6$ bits at the users and $N_{\text{L}}=18$ channel uses for exchanging $L_{1}$ and $L_{2}$ bits between the users. The coding rate is then $r_{1}=r_{2}=1/3$. For baselines, we first consider repetition coding, where User $i$ modulates each bit of ${\bf b}_{i}\in\{0,1\}^{L_{i}}$ with binary phase-shift keying (BPSK) and transmits each modulated symbol repetitively over $N_{\text{L}}/L_{i}=3$ channel uses. We consider tail-biting convolutional coding (TBCC) [29] as another baseline, adopted in LTE standards [30] for short block-length codes. We consider the trellis with $(7,[133,171,165])$ of a constraint length of $7$ and the generator $[133,171,165]$ in octal representation, and BPSK modulation. For our linear and RNN-based coding, there are several ways to exchange $L_{1}=L_{2}=6$ bits between the users over $N_{\text{L}}=18$ channel uses by varying the value of $K_{i}$. [^3] For linear coding, User $i$ modulates $L_{i}$ bits to $M_{i}=L_{i}/K_{i}$ messages with $2^{K_{i}}$-ary PAM, and $M_{i}$ messages are exchanged between the users. For RNN-based coding, we construct a coding architecture that supports $K_{1}$ and $K_{2}$ bits at the users. We consider the case of $K=K_{1}=K_{2}$ for both linear and RNN-based coding and investigate our coding schemes under different values of $K$.
@@
Figure 5: Sum-BLER with short block-lengths of $L_{1}=L_{2}=6$ and $N_{\text{L}}=18$ channel uses with $r_{1}=r_{2}=1/3$. RNN-based coding shows robustness to high channel noises, while linear coding performs well in very low noise scenarios. Here, sum-BLER of the channel coding schemes remains almost the same over $\text{SNR}^{\text{ch}}_{2}$, since sum-BLER is dominated by $\text{BLER}_{1}$ and $\text{BLER}_{1}$ is constant due to a constant value of $\text{SNR}^{\text{ch}}_{1}$. The proposed linear and RNN-based schemes benefit from interactive two-way coding and yield a significant performance improvement.
@@
Fig. 5 shows the sum-BLER performance under varying $\text{SNR}^{\text{ch}}_{2}$, where $\text{SNR}^{\text{ch}}_{1}=1$dB in Fig. 5(4(a)) and $\text{SNR}^{\text{ch}}_{1}=-1$dB in Fig. 5(4(b)). In Fig. 5(4(a)), the channel coding schemes, i.e., repetitive and TBCC, yield almost constant performance in sum-BLER, since sum-BLER is dominated by $\text{BLER}_{1}$ because of the fact that $\text{SNR}^{\text{ch}}_{1}<\text{SNR}^{\text{ch}}_{2}$, and $\text{BLER}_{1}$ is constant due to a constant value of $\text{SNR}^{\text{ch}}_{1}$. On the other hand, our two-way coding schemes, i.e., linear and RNN, yield better sum-BLER performance across different values of $\text{SNR}^{\text{ch}}_{2}$. The interactive exchange of information between the users allows User 2 under a better SNR condition to help User 1 under a worse SNR condition, which results in the performance improvement of $\text{BLER}_{1}$.
@@
For linear coding, we adopt the successive transmission for exchanging multiple messages between the users discussed in Sec. III-G, denoted by “Linear (ST)”. We also consider adopting an alternate channel use strategy for linear coding, denoted by “Linear”, which will be discussed in detail in the next subsection, Sec. VI-C. Note that this strategy improves the error performance as shown in the plots of Fig. 5, and thus we adopt this strategy throughout simulations. In Fig. 5(4(a)), linear coding with $K_{1}=K_{2}=3$ outperforms the other schemes when $\text{SNR}^{\text{ch}}_{2}$ is high, due to its benefit of utilizing more channel uses per message, which is allowed by high-order modulation discussed in Sec. III-G. Under the rest of the $\text{SNR}^{\text{ch}}_{2}$ regions, RNN-based coding yields the best performance, where we used two different coding architectures: (i) the softmax function in (67) and $K_{1}=K_{2}=6$ processing bits and (ii) the sigmoid function and $K_{1}=K_{2}=3$ bits. In Fig. 5(4(a)) with $\text{SNR}^{\text{ch}}_{1}=1$dB, RNN-based coding with the softmax function and $K_{1}=K_{2}=6$ bits performs better when $\text{SNR}^{\text{ch}}_{2}<10$dB, while RNN-based coding with the sigmoid function and $K_{1}=K_{2}=3$ bits performs better when $\text{SNR}^{\text{ch}}_{2}>10$dB. This implies that using the softmax function for decoding provides robustness to channel noises due to the block-level classification of the bit vector, which is aligned with the result of feedback-enabled GOWCs [17]. In Fig. 5(4(b)), where we consider a more noisy scenario with $\text{SNR}^{\text{ch}}_{1}=-1$dB, RNN-based coding with the softmax function yields the best performance over all the $\text{SNR}^{\text{ch}}_{2}$ regions, showing its robustness to the high channel noises.
VI-C Enhancement of the Linear Coding Scheme through Alternate Channel Uses
@@
In this subsection, we introduce an efficient strategy to further enhance our linear coding scheme through alternate channel uses. We consider that the users exchange $M=2$ message pairs, $(m_{1},m_{2})$ and $(m^{\prime}_{1},m^{\prime}_{2})$, over $N_{\text{L}}=6$ channel uses under power constraint $N_{\text{L}}P=6P$, where $\text{SNR}^{\text{ch}}_{1}=1$dB and $\text{SNR}^{\text{ch}}_{2}=20$dB. We first consider the successive transmission discussed in Sec. III-G. Each message pair is exchanged over $N=3$ channel uses under power constraint $3P$, where the corresponding encoding/decoding schemes are obtained via Algorithm 2. Fig. 6(5(a)) shows the power distribution for exchanging a message pair over $N=3$ channel uses with the proposed linear coding scheme. The two message pairs are then exchanged via the successive transmission with the power distribution in Fig. 6(5(b)). From the power distribution in Fig. 6(5(a)) and Fig. 6(5(b)), Users 1 and 2 only utilize about half of the channel uses, resulting in a waste of channel uses. More generally, when $N$ is odd, it has been shown that User 1 uses only odd-numbered channel uses, while User 2 uses only even-numbered channel uses and the last channel use. It interesting to see that although we do not impose any constraints on the separation of the channel usages between the two users, the separation of the channel usages is caused by solving the optimization problem. Specifically, the solution ${\bf x}$ of the sub-problem (48) has non-zero and zero values in alternating positions.
@@
Figure 6: Power distribution from linear coding along the channel uses. In (a), a pair of messages is exchanged over $N=3$ channel uses with power constraint $NP=3$. In (b), two pairs of messages are exchanged with power $N_{\text{L}}P=6$ over $N_{\text{L}}=6$ channel uses via the successive transmission of the two message pairs. Some of the channel uses are not utilized by the users, which is a waste of channel uses.
Motivated by the alternate pattern of the channel uses at the users, we introduce the strategy of alternate channel uses so that the two users fully utilize the channel uses for exchanging multiple pairs of messages. We consider the same example of exchanging $M=2$ message pairs, $(m_{1},m_{2})$ and $(m^{\prime}_{1},m^{\prime}_{2})$, over $N_{\text{L}}=6$ channel uses under power constraint $N_{\text{L}}P=6P$.
${\bf x}_{1}=[a_{1},0,b_{1},0,c_{1}],\quad{\bf x}_{2}=[0,a_{2},0,b_{2},c_{2}],$ (72)
Note that the transmit signals in (74) satisfy the power constraint $N_{\text{L}}P=6P$ while utilizing $N_{\text{L}}=6$ channel uses. This strategy also satisfies causality of the encoding process. We note that the transmit signal $c_{2}^{\prime}$ in $\bar{\bf x}_{2}$ of (74) is not restricted to follow a causal processing since it is only composed of a scaled version of the message $m_{2}^{\prime}$ from Proposition 2, and thus it can be transmitted at any channel use.
@@
Based on the alternate channel use strategy, we plot the power distribution in Fig. 7. Fig. 7(6(a)) shows the power distribution for exchanging a pair of messages over $N=5$ channel uses discussed in Step 1, while Fig. 7(6(b)) shows the power distribution for exchanging two pairs of messages over $N_{\text{L}}=6$ channel uses when adopting the alternate channel use strategy discussed in Step 2. This alternate channel use strategy allows the users to fully utilize the channel uses within the same power constraint compared to the successive transmission in Fig. 6(5(b)), and thus leads to the error performance improvement as shown in Fig. 5. Throughout simulations, we adopt this strategy for linear coding.

### v2

@@ V Computational Complexity of Linear and RNN-Based Coding
In this section, we analyze the computational complexity of our proposed linear and RNN-based coding schemes. We first look into the case of linear coding. In Algorithm 2, we solve for the encoding schemes, i.e., ${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, and ${\bf F}_{2}$, given the fixed number of channel uses, $N$, and the noise variances, $\sigma_{1}^{2}$ and $\sigma_{2}^{2}$. Subsequently, we obtain the decoding (combining) schemes, ${\bf w}_{1}$ and ${\bf w}_{2}$, by (35) and (25) as functions of the encoding schemes. Once the encoding/decoding schemes are determined, they are used for encoding via (10)-(11) and decoding described in Sec. III-B without being recalculated. Specifically, the complexity of encoding at User 1 is $\mathcal{O}(N^{2})$ from (10), where ${\bf F}_{1}{\bf F}_{2}$ can be pre-calculated, while User 2 has the complexity of $\mathcal{O}(N)$ in (11), where the computation of ${\bf F}_{2}{\bf y}_{2}$ has $\mathcal{O}(N)$ complexity since ${\bf F}_{2}$ is in the form of (47). For decoding, User 1 has complexity of $\mathcal{O}(N^{2})$, where the decoding process at User 1 consists of (i) $\tilde{\bf y}_{1}=({\bf I}+{\bf F}_{2}{\bf F}_{1})^{-1}({\bf y}_{1}-{\bf F}_{2}{\bf g}_{1}m_{1})$ and (ii) ${\hat{m}}_{2}={\bf w}_{2}^{\top}\tilde{\bf y}_{1}$, described above (28). Here, $({\bf I}+{\bf F}_{2}{\bf F}_{1})^{-1}$ and ${\bf F}_{2}{\bf g}_{1}$ can be pre-computed. On the other hand, User 2 has $\mathcal{O}(N)$ complexity, where the decoding process at User 2 is composed of (i) $\tilde{\bf y}_{2}={\bf y}_{2}-{\bf F}_{1}{\bf g}_{2}m_{2}$ and (ii) ${\hat{m}}_{1}={\bf w}_{1}^{\top}\tilde{\bf y}_{2}$, described above (38). Here, ${\bf F}_{1}{\bf g}_{2}$ can be pre-computed.
@@
The encoding and decoding complexity at the users for linear coding and RNN-based coding are summarized in Tables II and II. Linear coding has a low coding complexity due to the simplified system model adopting linear operation for encoding and decoding in (10)-(11), (25), and (35). While the RNN-based coding causes higher encoding/decoding complexity, it benefits from higher degrees of flexibility thanks to the non-linearity provided by deep neural networks, leading to better error performance under many practical noise scenarios (shown in Sec. VI).
@@ VI Experimental Results
In this section, we investigate both our linear and learning-based coding methodologies through numerical experiments. We first describe the simulation setup (Sec. VI-A) and validate our coding methodologies in a short block-length regime (Sec. VI-B). In addition, we discuss the enhancement of our linear coding scheme by thoroughly investigating its solution behavior (Sec. VI-C). We also show the effectiveness of our coding methodologies in a medium/long block-length regime (Sec. VI-D). Then, we analyze the benefits of two-way coding (Sec. VI-E), and show that the best coding strategy is different in SNR regimes (Sec. VI-F). Moreover, we investigate the error performance by varying the rates (Sec. VI-G) and provide information theoretic insights on power distribution of both linear and RNN-based coding (Sec. VI-H). Finally, we analyze the block-length gain of our RNN-based coding (Sec. VI-I).
@@ VI-A Simulation Parameters
For RNN-based coding, the number of training data samples that we used is $J=10^{7}$, the batch size is $N_{\text{batch}}=2.5\times 10^{4}$, and the number of epochs is $N_{\text{epoch}}=100$. We use the ADAM optimizer [27] and a decaying learning rate, where the initial rate is 0.01 and the decaying ratio is $\gamma=0.95$ applied for every epoch. We also use gradient clipping for training, where the gradients are clipped when the norm of gradients is larger than 1. We adopt two layers of uni-directional GRUs at the encoder and two layers of bi-directional GRUs at the decoder for each user, with $N_{\text{neurons}}=N_{i}^{(\text{enc})}=N_{i}^{(\text{dec})}=50$ neurons at each GRU. We initialize each neuron in a GRU with the uniform distribution $U(-\zeta,\zeta)$, where $\zeta=1/\sqrt{N_{\text{neurons}}}$, and all the power weights and attention weights to 1. We train our neural network model under particular noise scenarios, i.e., $\text{SNR}_{1}^{\text{ch}}$ and $\text{SNR}_{2}^{\text{ch}}$, and conduct inference under the same noise scenario. For inference, we consider $10^{10}$ data tuples to calculate BER and BLER.
@@ VI-B Short Block-Length
[…]
@@
Figure 5: Sum-BLER with short block-lengths of $L_{1}=L_{2}=6$ bits and $N_{\text{L}}=18$ channel uses with $r_{1}=r_{2}=1/3$. RNN-based coding shows robustness to high channel noises, while linear coding performs well in very low noise scenarios. Here, sum-BLER of the channel coding schemes remains almost the same over $\text{SNR}^{\text{ch}}_{2}$, since sum-BLER is dominated by $\text{BLER}_{1}$ and $\text{BLER}_{1}$ is constant due to a constant value of $\text{SNR}^{\text{ch}}_{1}$. The proposed linear and RNN-based schemes benefit from interactive two-way coding and yield a significant performance improvement.
@@
Fig. 5 shows the sum-BLER performance under varying $\text{SNR}^{\text{ch}}_{2}$, where $\text{SNR}^{\text{ch}}_{1}=1$dB in Fig. 5(5(a)) and $\text{SNR}^{\text{ch}}_{1}=-1$dB in Fig. 5(5(b)). The channel coding schemes, i.e., repetitive and TBCC, and the benchmark Sum-OL-LB yield almost constant value in sum-BLER when $\text{SNR}^{\text{ch}}_{1}<\text{SNR}^{\text{ch}}_{2}$, since sum-BLER is dominated by $\text{BLER}_{1}$. Note that $\text{BLER}_{1}$ is constant due to a constant value of $\text{SNR}^{\text{ch}}_{1}$. On the other hand, our two-way coding schemes, i.e., linear and RNN, yield better sum-BLER performance across different values of $\text{SNR}^{\text{ch}}_{2}$. The interactive exchange of information between the users allows User 2, which has a better channel SNR than that of User 1, to help User 1, which results in a performance improvement of $\text{BLER}_{1}$, ultimately improving sum-BLER. It is interesting that cooperation enables users to achieve a lower sum-BLER than Sum-OL-LB. This is because User 1’s error performance is improved by the feedback from User 2, resulting in an error rate even lower than the open-loop lower bound. The behavior of individual BLER curves is further analyzed in Sec. VI-E and Fig. 9. This phenomenon is also observed in feedback-enabled one-way communications; as shown in Fig. 2 in [15], using feedback in coding improves the BLERs, achieving performance better than the open-loop lower bound.
@@
For linear coding, we adopt the successive transmission for exchanging multiple messages between the users discussed in Sec. III-G, denoted by “Linear (ST)”. We also consider adopting the alternate channel use strategy for linear coding, denoted by “Linear”, discussed in Sec. III-H. Note that this strategy improves the error performance as shown in the plots of Fig. 5, and thus we adopt this strategy throughout simulations. As discussed in Sec. III-G, the value of $K_{i}$ for linear coding affects sum-error performance in asymmetric channels due to the trade-off between the minimum distance of the constellation and the feedback benefit depending on $K_{i}$. In Figs. 5(5(a)) and (5(b)), at high channel SNR at User 2, e.g., at $\text{SNR}^{\text{ch}}_{2}=30$dB, high modulation up to $K_{i}=3$ gives better performance since the feedback benefit from using a larger number of channel uses $N=K_{i}/r_{i}$ outweighs the performance loss due to the high modulation order $2^{K_{i}}$ and the subsequently decreased constellation distance. Under lower channel SNRs, $K_{i}=2$ performs the best at $\text{SNR}^{\text{ch}}_{2}=20$dB, while $K_{i}=1$ yields the best performance when $\text{SNR}^{\text{ch}}_{2}\leq 15$dB. This indicates that, at lower channel SNRs, smaller values of $K_{i}$ perform better than larger values. This is because, at low channel SNRs, higher modulation order results in a significant loss in error performance due to the decreased constellation distance, which outweighs the benefits of feedback. Note that with $K_{i}=6$, the sum-error performance increases drastically across all User 2’s channel SNR values since the performance loss due to the high modulation order dominates over the feedback benefit. In Fig. 5(5(a)), we observe that linear coding with $K_{1}=K_{2}=3$ outperforms the other schemes when $\text{SNR}^{\text{ch}}_{2}$ is high. For the rest of the $\text{SNR}^{\text{ch}}_{2}$ regions, RNN-based coding yields the best performance. In Fig. 5(5(a)) with $\text{SNR}^{\text{ch}}_{1}=1$dB, RNN-based coding with the softmax function and $K_{1}=K_{2}=6$ bits performs better when $\text{SNR}^{\text{ch}}_{2}<10$dB, while RNN-based coding with the sigmoid function and $K_{1}=K_{2}=3$ bits performs better when $\text{SNR}^{\text{ch}}_{2}>10$dB. This implies that using the softmax function for decoding provides robustness to channel noises due to the block-level classification of the bit vector, which is aligned with the result of feedback-enabled GOWCs [18]. In Fig. 5(5(b)), where we consider a more noisy scenario with $\text{SNR}^{\text{ch}}_{1}=-1$dB, RNN-based coding with the softmax function yields the best performance over all the $\text{SNR}^{\text{ch}}_{2}$ regions, showing its robustness to the high channel noises.
VI-C Linear Coding Enhanced by Alternate Channel Uses
@@
We consider that the users exchange $M=2$ message pairs, $(m_{1},m_{2})$ and $(m^{\prime}_{1},m^{\prime}_{2})$, over $N_{\text{L}}=6$ channel uses under power constraint $N_{\text{L}}P=6P$, where $\text{SNR}^{\text{ch}}_{1}=1$dB and $\text{SNR}^{\text{ch}}_{2}=20$dB. We first consider the successive transmission discussed in Sec. III-G. Fig. 6(6(a)) shows the power distribution for exchanging a message pair over $N=3$ channel uses with the proposed linear coding scheme. The two message pairs are then exchanged via the successive transmission with the power distribution in Fig. 6(6(b)). From the power distribution in Fig. 6(6(a)) and Fig. 6(6(b)), Users 1 and 2 only utilize about half of the channel uses, resulting in a waste of channel uses. Next, we consider the same setup for the strategy of alternate channel uses. Based on the alternate channel use strategy, we plot the power distribution in Fig. 7. Fig. 7(7(a)) shows the power distribution for exchanging a pair of messages over $N=5$ channel uses discussed in Step 1 in Sec. III-H. Fig. 7(7(b)) shows the power distribution for exchanging two pairs of messages over $N_{\text{L}}=6$ channel uses when adopting the alternate channel use strategy discussed in Step 2 in Sec. III-H. This alternate channel use strategy allows the users to fully utilize the channel uses within the same power constraint compared to the successive transmission in Fig. 6(6(b)), and thus leads to the error performance improvement as demonstrated in Fig. 5. Throughout simulations, we adopt this strategy for linear coding.
@@
Figure 6: Power distribution from linear coding along the channel uses. In (a), a pair of messages is exchanged over $N=3$ channel uses with power constraint $NP=3$. In (b), two pairs of messages are exchanged with power $N_{\text{L}}P=6$ over $N_{\text{L}}=6$ channel uses via the successive transmission of the two message pairs. Some of the channel uses are not utilized by the users, which causes a waste of channel uses.